OpenStudy (trojanpoem):

Solve the O.D.E: y^2 y'' + 4 = 0

OpenStudy (faiqraees):

Maybe it can be considered in the form \[ y''+0y'+\frac{4}{y^2}=0\]

OpenStudy (trojanpoem):

I thought of doing the substitution which yields a first order O.D.E: y' = P y'' = p dp/dy y^2 p dp/dy + 4 = 0 pdp = -4/y^2 p^2/2 = 4/y + C p^2 = 8/y + C p = sqrt( 8/y + c) dy/dx = sqrt(8/y + c) can't seem to find a way to integrate dy/sqrt(8/y + c)

OpenStudy (trojanpoem):

y'' + 0 y' + 4/y^2 = 0 Ok, then ?

OpenStudy (faiqraees):

And then, I am not sure thought, we can treat it whole as a second order linear differential equation and can use the general solutions to find its solution

OpenStudy (trojanpoem):

Differential operator ? variation of parameters ? undetermined coefficient ?

OpenStudy (trojanpoem):

y'' = -4/y^2 D^2 y = -4 /y^2 D^2 y = -4 /y^2 To me stalemate ..

OpenStudy (faiqraees):

Oh no no. Havent you studied the general solutions for some simple differential equations like y"+y+a =Ae^cx +Be^dx etc?

OpenStudy (trojanpoem):

brah , x ? not y

OpenStudy (trojanpoem):

y'' + 0 y' + 4/y^2 = 0 y'' + y + a = ... because you let y = e^bx y' = be^bx y'' = b^2 e^bx

OpenStudy (trojanpoem):

And BTW, what you have given is called a complementary function

OpenStudy (trojanpoem):

The general solution is a combination of both a complementary function and a particular solution to the equation.

OpenStudy (faiqraees):

Yes Its been a while since I have done these so...

OpenStudy (faiqraees):

So you know those right?

OpenStudy (faiqraees):

Because this is what I thought to be solved Maybe it can be considered in the form \[y''+0y'+1=1-\frac{4}{y^2}\]

OpenStudy (faiqraees):

Now you have both complimentary function and particular Integral. @IrishBoy123 Maybe you can help

OpenStudy (irishboy123):

this is non-linear, right?

OpenStudy (trojanpoem):

True.

OpenStudy (irishboy123):

so old trick, sorry i have not read the therad, is \(y^2 \dfrac{d^2 y}{dx^2} = - 4\) \( y^2 y' \dfrac{d y'}{dy} = -4\)

OpenStudy (trojanpoem):

I did it :P , got: dy/sqrt(8/y + c) = x + b

OpenStudy (irishboy123):

i'd ask wolfram, personally you're changing variable here....

OpenStudy (trojanpoem):

I did so in the second comment. Actually the book final answer is cute. 4 + ay^2 = b^2(x+b)^2

OpenStudy (trojanpoem):

.

OpenStudy (irishboy123):

that looks do-able are you struggling to get it, or just sharing the answer?

OpenStudy (trojanpoem):

Struggling to get it.

OpenStudy (holsteremission):

As a first step, \[y^2\frac{\mathrm d^2y}{\mathrm dx^2}+4=0\implies \frac{\mathrm d^2y}{\mathrm dx^2}=-\frac{4}{y^2}\implies \frac{\mathrm dy}{\mathrm dx}\frac{\mathrm d^2y}{\mathrm dx^2}=-\frac{4}{y^2}\frac{\mathrm dy}{\mathrm dx}\]Integrating both sides with respect to \(x\) yields \[\int\frac{\mathrm dy}{\mathrm dx}\frac{\mathrm d^2y}{\mathrm dx^2}\,\mathrm dx=\frac{1}{2}\int\frac{\mathrm dy}{\mathrm dx}\,\mathrm d\left(\frac{\mathrm dy}{\mathrm dx}\right)=\frac{1}{2}\left(\frac{\mathrm dy}{\mathrm dx}\right)^2+C_1\]on the left, while on the right you get \[-4\int\frac{1}{y^2}\frac{\mathrm dy}{\mathrm dx}\,\mathrm dx=-4\int\frac{\mathrm dy}{y^2}=\frac{4}{y}+C_2\]so you're left with \[\frac{1}{2}(y')^2=\frac{4}{y}+C\]which appears a bit easier to work with.

OpenStudy (trojanpoem):

Same as the substitution , of P or y' yields the integration of dy/sqrt( 8/y + C) I find it cool to use dy/dx confusing xD

OpenStudy (irishboy123):

\(y'=\sqrt { \frac{8}{y}+C}\) \(\int \dfrac{1}{\sqrt { \frac{8}{y}+C}}dy=\int dx\) \(\int \dfrac{y}{\sqrt { 8+C y^2}}dy= \int dx\) and pattern match from \(\dfrac{d}{dx} \sqrt { 8+C y^2} = \dfrac{1/2 * 2 Cy}{\sqrt { 8+C y^2}}\) to solve the LHS

OpenStudy (irishboy123):

i missed out a \(\pm\) in there but if you square it as they do, that problem goes away

OpenStudy (trojanpoem):

I think you missed a y ? sqrt(8/y + c) = y/y * sqrt(8/y + c) = sqrt(8y + cy^2) ?

OpenStudy (trojanpoem):

sudden non-scheduled windows upgrade to version 1607 Lagging as hell. PC shutdown automatically

OpenStudy (trojanpoem):

@IrishBoy123

OpenStudy (irishboy123):

@TrojanPoem you're right! :-( so yes, that was terrible algebra on my part

OpenStudy (trojanpoem):

Stuck with two more ... I manged to get to a form integrate-able by a trigonometric substitution. It can never be simplified to the book's ANS.